8.00 L of argon is allowed to expand at 21.0 ° C to a volume of 25.0 L. If the pressure remains constant, what is the final temperature (in degrees Celsius)?


1
Expert's answer
2021-05-18T07:52:20-0400

We have two conditions:


First condition:

"V_1=8l"

"T_1=21^0 C"


Second condition:

"V_2=25l"

"letT_2=T_2"


As pressure ,moles ,R is constant then

"V\propto T"


Hence:


"\dfrac{V_1}{V_2}=\dfrac{T_1}{T_2}"


"\dfrac{8}{25}=\dfrac{21}{T_2}"


"T_2= \dfrac{21\times25}{8}=65.625^0 C"